OpenStudy (anonymous):

HELP ME PLEASE! Typical exponential function, y=a^x has an asymptote ____ and y-intercept ____ . y=0; (0,a) y=0; (0,1) y=a; (0,a) y=a; (0,1)

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

the intercept is 1, right ?

OpenStudy (accessdenied):

That is correct, x=0; a^0 = 1. :)

OpenStudy (anonymous):

How do I find the asymptote ?

OpenStudy (accessdenied):

Consider what happens when you use very very large values of x in both the positive and negative directions. a^x will explode when you go for huge numbers in positive direction. However, negative values: a^(-x) = 1/a^(x), this part goes towards zero because for larger and larger x-values, we get 1/huge number.

OpenStudy (anonymous):

Yes...

OpenStudy (accessdenied):

That case is for a>1, I should add. For a<1 the opposite is true, but it still goes towards y=0 in the other direction. You could plot a few examples: y = 2^x, y = 1/2^x, etc. to get a good idea of their general form. :).

OpenStudy (anonymous):

So the asymptote is a?

OpenStudy (anonymous):

I don't get how does it approaches zero, if you plug in bigger numbers?

OpenStudy (accessdenied):

You know the asymptote means the value it tends towards for the extremes? 

OpenStudy (anonymous):

that I did NOT know-:( Thank you